# Exposía: Academic Writing Assessment of Exposés and Peer Feedback

### v1

@@ I.2 Review Scoring Prompt
Appendix K Runtime and Token Consumption
@@
Tables 24 and 25 report runtime and token usage for exposé and review scoring, respectively. Across both tasks, combined prompting is essential for throughput: it requires a single forward pass per submission and therefore yields the lowest latency (110–430s on average across models for exposé scoring). In contrast, single-criterion prompting multiplies calls per submission, increasing total runtime to 284–1133s on average for exposés, with worst cases up to 23 minutes. The same multiplicative effect appears in token processing: although single-criterion outputs are shorter per call, repeating the full input across criteria yields an order-of-magnitude increase in total processed tokens per submission (approximately $13\times$ in our setup). Review scoring exhibits the same trend (Table 25), with combined prompting consistently yielding lower latency and substantially fewer processed tokens per submission than single-criterion prompting.
@@
Appendix L Behaviour Data: Interaction Logs and Semester Dynamics

### v2

@@ Appendix I Experimental Details
Figure 7: Exposé: Criterion $\times$ model heatmap (JSD). Cell values show weighted mean JSD per criterion and model (weights: items_in_bin). Lower is better. Model abbreviations: Opus-4.6=Claude Opus 4.6; Gem-3.1P=Gemini 3.1 Pro; GPT-5.2=GPT 5.2; OSS-120=GPT OSS (120B); OSS-20=GPT OSS (20B); L3.3-70=Llama 3.3 (70B); Qwen-80=Qwen 3 (80B).
Figure 8: Reviews: Criterion $\times$ model heatmap (JSD). Cell values show weighted mean JSD per criterion and model (weights: items_in_bin). Lower is better. Model abbreviations: Opus-4.6=Claude Opus 4.6; Gem-3.1P=Gemini 3.1 Pro; GPT-5.2=GPT 5.2; OSS-120=GPT OSS (120B); OSS-20=GPT OSS (20B); L3.3-70=Llama 3.3 (70B); Qwen-80=Qwen 3 (80B).
Appendix L Runtime and Token Consumption
@@
Tables 26 and 27 report runtime and token usage for exposé and review scoring, respectively. Across both tasks, combined prompting is essential for throughput: it requires a single forward pass per submission and therefore yields the lowest latency (110–430s on average across models for exposé scoring). In contrast, single-criterion prompting multiplies calls per submission, increasing total runtime to 284–1133s on average for exposés, with worst cases up to 23 minutes. The same multiplicative effect appears in token processing: although single-criterion outputs are shorter per call, repeating the full input across criteria yields an order-of-magnitude increase in total processed tokens per submission (approximately $13\times$ in our setup). Review scoring exhibits the same trend (Table 27), with combined prompting consistently yielding lower latency and substantially fewer processed tokens per submission than single-criterion prompting.
@@
Appendix M Behaviour Data: Interaction Logs and Semester Dynamics